OpenStudy (anonymous):

Show the expanded form and simplification of x^6 over x^2 . Explain in your own words how you can simplify X^6 over x^2 without having to write the expanded form. Part 2: Create your own fraction with like bases, coefficients, and show its simplification.

OpenStudy (anonymous):

you know that x divided by x is equal to 1 here you have "six times x" divided by "two times x": x.x.x.x.x.x/x.x x/x=1 then you get 1.1.x.x.x.x = x^4 in other words you can say that x^6/x^2 = x^(6-2) = x^4